OpenStudy (anonymous):

(In the form of 0*infinity) Find the limit as x approaches pi/2 from the negative of (pi/2-x)secx

OpenStudy (phi):

write it as \[ \frac{ \frac{\pi}{2} - x}{\cos(x)} \] now it is in the form 0/0